### v1

# Geodetic Exploration of the Interior Structure of Enceladus

Abstract
Enceladus is among the most intriguing bodies in the solar system due to its astrobiological potential. Determining the extent and duration of sustainable habitability requires characterizing the interior properties, and in particular, the extent and distribution of the tidal heating.The intensity of any inferred geophysical activity in the core has direct implications for understanding the hydrothermal activity and the budget of biologically useful species. We build a statistical framework using an MCMC approach to constrain the interior based on currently available observations of libration, shape, heat budget, and mass. We use this framework to examine the extent that geodetic measurements can constrain the interior structure, with an emphasis on determining the partitioning of tidal dissipation between the shell and the core. We quantify plausible ranges of gravitational and displacement tidal Love numbers consistent with existing observations. We demonstrate that measuring the amplitude and phase of $k_{2}$ alone can only constrain the total tidally dissipated energy, without a sensitivity to the distribution of heat. However, measuring the amplitude and phase of $h_{2}$ or $l_{2}$ facilitates determination of the shell rheology. We provide the precisions required for measuring $k_{2}$, $h_{2}$, and/or $l_{2}$ that enable distinguishing between the two main scenarios of tidal heating, i.e., in the shell versus the core. We also investigate the effect of the structural heterogeneities of the shell on the tidal response. Lastly, we evaluate the efficacy of a suite of future geodetic measurements to constrain key interior properties essential to instantaneous and sustained habitability at Enceladus.
Tides, Tidal heating, Love numbers, Libration, Gravity, Icy moons
show]abagheri@caltech.edu
1 Introduction and Background
Flybys of Saturn’s moon Enceladus by the Cassini spacecraft revealed multiple intriguing features suggesting an active surface and interior: plumes that continuously erupt from a region approximately centered on the satellite’s south pole (Porco et al., 2006) which feed Saturn’s diffuse E-ring with water-ice crystals (Kempf et al., 2010); a very bright surface (Hendrix et al., 2010); a high surface heat flux; and, importantly, evidence for a subsurface ocean (Porco et al., 2006; Spencer et al., 2006; Howett et al., 2011; McKinnon, 2015; Thomas et al., 2016; Park et al., 2024). This subsurface water body likely meets the criteria for instantaneous habitability (Cockell et al., 2024): An extensive liquid water reservoir containing salts and complex organic compounds and the expected chemical disequilibrium required for supporting metabolism. These conditions can occur through water-rock interactions, presumably in inferred hydrothermal systems at or below the seafloor (Postberg et al., 2009a; Howett et al., 2011; Glein et al., 2015; Hsu et al., 2015; Cable et al., 2021; Cockell et al., 2024). These lines of evidence for a potentially habitable environment position Enceladus as a compelling target for future exploration by orbiters and landers (NASEM, 2023). Studying to what extent these requirements have been sustained at Enceladus requires understanding the geophysical processes in the interior over geological timescales (see the discussions of sustained or dynamical habitability in NASEM (2023); Cockell et al. (2024)).
Analysis of Cassini’s imaging, thermal mapping, and Deep Space Network (DSN) radiometric tracking data explored the global shape, gravity field, libration (i.e., diurnal wobble of the ice shell), and the surface topography of Enceladus, providing insights into the interior layers’ properties. The large amplitude of forced physical libration suggests that the ice crust is mechanically decoupled from the rocky core, likely by a global ocean (Nadezhdina et al., 2016; Thomas et al., 2016; Van Hoolst et al., 2016; Park et al., 2024). The presence of a global ocean is also supported by analysis of the low-order gravity field (McKinnon, 2015; Čadek et al., 2016). Estimates of the ice shell thickness and the surface temperatures require a total heat loss of approximately 25-40 GW (Hemingway & Mittal, 2019; Nimmo et al., 2023), while the measured heat output from the South Polar Terrain (SPT) alone is 4-19 GW (Howett et al., 2011). With an estimated radiogenic heating rate of less than 0.3 GW (Vance et al., 2007; Roberts & Nimmo, 2008) and minimal additional heating due to libration (Rambaux et al., 2010; Shao & Nimmo, 2022), the most likely heat source is therefore dissipation due to the eccentricity tides exerted by Saturn (e.g., Nimmo et al., 2023). Initial analyses suggested that given the global properties of Enceladus and Saturn, tidal dissipation is insufficient to explain the estimated heat loss (Meyer & Wisdom, 2007). However, later studies demonstrated that Saturn is supplying tidal energy and angular momentum to its satellites at a higher rate than previous estimates (Lainey et al., 2017), due to resonance locking between the moons and the planet’s internal oscillation modes (Fuller et al., 2016; Lainey et al., 2017, 2020; Jacobson, 2022).
It remains unknown whether tidal dissipation primarily occurs in Enceladus’s core or ice shell. Distinguishing between the existing end-member scenarios is vital for assessing Enceladus’s sustained habitability (e.g., NASEM, 2023; Cockell et al., 2024). A tidally active core can facilitate hydrothermal activity at the bottom of the ocean and the continuous contact between water and fresh rock, potentially supporting the chemical reactions necessary for life to develop and evolve over geological timescales (Vance et al., 2007; Hsu et al., 2015; Choblet et al., 2017). Therefore, a comprehensive description of the global heat budget allows the distinction between a primordial and a present-day hydrothermal source of organics, volatiles, and nanosilica materials in the plumes (Vance et al., 2007; Hsu et al., 2015; Choblet et al., 2017). Studies have considered viscous dissipation in the ice shell as the mechanism of maintaining the global ocean (Roberts & Nimmo, 2008; Shoji et al., 2013; Běhounková et al., 2015; Efroimsky, 2018; Kang & Flierl, 2020). These models require a mean viscosity of $\eta_{shell}\approx 10^{13-15}$ Pa.s which is lower than the estimated ice viscosity at the temperatures in Enceladus. Alternative studies propose dissipation in water-filled fractures (Kite & Rubin, 2016), in the ocean (Tyler, 2020), or in a highly porous and dissipative rocky core with a very low effective viscosity (Roberts, 2015; Choblet et al., 2017; Neumann & Kruse, 2019; Souček et al., 2019; Liao et al., 2020; Aygün & Čadek, 2023).
Considering the ambiguous understanding of the interior structure, geophysical investigations of Enceladus must account for the possibility of heating in the deep interior. Ermakov et al. (2021) and Genova et al. (2024) propose an effective approach to explore the interior by measuring the potential tidal Love number $k_{2}$ along with the libration or gravity-topography admittance to constrain the interior structure (i.e., thicknesses and densities of layers) at Enceladus. However, the ranges of interior parameters considered in previous studies inherently assume negligible tidal dissipation in the core without providing a quantitative approach to discriminating between the different scenarios of tidal dissipation. As we show later, a tidally active core will affect the time-dependent gravity field of the body and produce a gravitational signal which resembles that of a tidally dissipative ice shell, resulting in a non-unique interpretation of the amplitude and phase of $k_{2}$. Also, gravity and topography measurements have revealed structural heterogeneities in the Enceladus ice shell (Yin & Pappalardo, 2015; Park et al., 2024; Schenk & McKinnon, 2024). Deviations of the body from spherical symmetry can potentially affect its tidal response (e.g., Dmitrovskii et al., 2022; Berne et al., 2023; Goossens & Sabaka, 2025) by driving deformation at higher harmonic degrees (e.g., $l=3,4$), and mode coupling which is expected to produce variations in degree-2 Love numbers evaluated at different spherical harmonic orders, i.e., $k_{2m}$ and $h_{2m}$ (e.g., Bagheri et al., 2021; Pou et al., 2022; Berne et al., 2023).
Here, we revisit the range of the interior model parameters based on the currently available observations using a MCMC statistical framework. We use this framework to examine the effectiveness of a suite of geodetic measurements including static gravity field, complex tidal Love numbers, libration, and heat loss rate that can constrain the interior of Enceladus. We highlight the degeneracy in interpreting $k_{2}$ to locate the tidally active region of Enceladus and provide a roadmap to distinguish between the two classes of tidal dissipation. We use our framework to compute the required precisions for measuring the tidal and displacement Love numbers. Using the proposed suite of future geophysical measurements and synthetic observations that can become available in future missions, we consider constraining interior parameters, such as the thickness, density, and rheology of the interior layers. These constraints can significantly improve our understanding of the instantaneous and sustained habitability at Enceladus.
2 Methods
We describe the geophysical models needed to compute the geodetic quantities in our forward model. The models include the tidal response in the form of the complex Love numbers and maximum surface deformations, the viscoelastic models used in computing the tidal response and physical libration, and calculation of the static gravity coefficients. Detailed derivations are provided in the Appendix.
2.1 Tidal Response
The eccentric orbit of Enceladus around Saturn drives periodic gravitational tidal forcing of the moon. This forcing potential, truncated at spherical harmonic degree 2 and to first order for eccentricity, can be written as (Murray & Dermott, 1999):
$U_{2}(r,\theta,\phi,t)=r^{2}\omega^{2}e\big{[}-\frac{3}{2}P_{2}^{0}(\lambda)\cos\omega t+P_{2}^{2}(\lambda)(\frac{3}{4}\cos\omega t\cos 2\phi+\sin\omega t\sin 2\phi)\big{]},$ (1)
where $r$ is the radius, $\theta$ is the co-latitude, $\phi$ is the longitude, $t$ is time, $\omega$ is the frequency of the forcing tide equal to the orbital frequency 53.07 $\mu\rm Rad.s^{-1}$, $e$ is the orbital eccentricity equal to 0.0047, $P_{2}^{0}$ and $P_{2}^{2}$, are the un-normalized associated Legendre polynomials, and $\lambda=\cos\theta$. For a spherically symmetric body, the tidal deformations in spherical coordinates in radial and tangential directions are described by the displacement Love numbers $h_{2}$ and $l_{2}$ (Melchior, 1983):
$u_{r}(r,\theta,\phi,t)=\frac{h_{2}(r)}{g(r)}U_{2}(r,\theta,\phi,t),~{}$
$u_{\theta}(r,\theta,\phi,t)=\frac{l_{2}(r)}{g(r)}\frac{\partial U_{2}(r,\theta,\phi,t)}{\partial\theta},~{}$
$u_{\phi}(r,\theta,\phi,t)=\frac{1}{\sin\theta}\frac{l_{2}(r)}{g(r)}\frac{\partial U_{2}(r,\theta,\phi,t)}{\partial\phi},$ (2)
where $g(r)$ is the gravitational acceleration at radius $r$. The resulting perturbation in the gravitational potential field of Enceladus ($W_{2}$) is calculated through gravitational tidal Love number $k_{2}$ as:
$W_{2}(r,\theta,\phi,t)=k_{2}(r)U_{2}(r,\theta,\phi,t).$ (3)
$k_{2}$, $h_{2}$, and $l_{2}$ are sensitive to the properties of the entire body and primarily depend on the density and rheological structure. The maximum displacements in the radial and horizontal directions are expressed as (Park et al., 2020):
$\max(\Delta R)=12\sqrt{2/7}\frac{|h_{2}|R_{E}^{2}\omega^{2}e}{g},$
$\max(\Delta E)=\frac{12|l_{2}|R_{E}^{2}\omega^{2}e}{g},$
$\max(\Delta N)=\frac{9|l_{2}|R_{E}^{2}\omega^{2}e}{g},$ (4)
where $R_{E}$ is the radius of Enceladus, and $\Delta R$, $\Delta E$, and $\Delta N$ are the radial, easterly, and northerly displacements on the surface, respectively. Because of the viscoelastic behavior of materials, the tidal Love numbers are complex values, implying that there is a phase lag between the forcing potential and the tidal response. We compute the total heat generated by tides from the complex $k_{2}$ (Segatz et al., 1988):
$\dot{E}=\frac{21}{2}\frac{n^{5}R_{E}^{5}}{G}e^{2}\frac{|k_{2}|}{Q},$ (5)
where $\dot{E}$ is the total power generated by tides, $n$ is the orbital mean motion, $G$ is the gravitational constant, and $Q$ is the tidal quality factor given by (Goldreich & Soter, 1966):
$Q^{-1}=\sin(|\epsilon_{k_{2}}|).$ (6)
Here, $\epsilon_{k_{2}}$ is the phase lag of the perturbation in the gravitational potential, which can represent the total tidal energy generated throughout the body. In most cases, the tidal quality factor $Q$ is simply defined as $\tan(\epsilon_{k_{2}})$. However, this simplification is only correct when the phase lag is small such that $\tan(\epsilon_{k_{2}})\approx\sin(\epsilon_{k_{2}})$. For bodies like Enceladus and other satellites that may experience intense tidal heating and hence high tidal phase lag, this approximation is no longer valid (Efroimsky, 2012a, b). Here, we explicitly consider phase lags, as opposed to the quality factor $Q$. This choice is appropriate since for Enceladus the phase lags are potentially large, $Q$ is defined inconsistently in the literature, and our focus is on direct observables. We also generally assume that the total heat loss is equal to the tidally generated heat computed by Equation (5); in other words, we assume that Enceladus is currently in a thermal steady state.
2.2 Viscoelastic model
In order to predict the tidal response of Enceladus, we must consider rheological models that incorporate shear dissipation in rock and icy regions. When both viscous and elastic mechanisms are active, the viscoelastic response depends on the shear modulus $\mu$, frequency of excitation $\omega$, and the effective viscosity $\eta$. The Maxwell viscoelastic model provides a reasonable description of a material’s behavior at time scales that are significantly shorter or longer than the Maxwell (relaxation) time of the body defined as $\tau_{M}=\eta/\mu$. $\tau_{M}$ represents the approximate timescale at which the behavior transitions from being dominated by the elastic regime to the viscous regime. However, a Maxwell model may not be appropriate for describing tidal dissipation in planetary bodies that have tidal periods ranging from hours to a few days, which may be near the relaxation time of the body. Thus, applicability of Maxwell’s model for capturing tidal dissipation has been questioned for multiple planetary bodies (Bills et al., 2005; Castillo-Rogez et al., 2011; Steinbrügge et al., 2018; Bagheri et al., 2019; Petricca et al., 2024). Several models based on micro-scale physical processes and laboratory measurements have been proposed to mimic the behavior of a material at time scales between the elastic and viscous regimes (e.g., Castillo-Rogez et al., 2011; Bagheri et al., 2019, 2021; Petricca et al., 2024; Tobie et al., 2025). In particular, the Andrade model has been recommended to model tidal dissipation in celestial bodies taking into account the effect of anelastic behavior in the transition from dominantly elastic to dominantly viscous regimes (Castillo-Rogez et al., 2011; Renaud & Henning, 2018; Padovan et al., 2014; Bagheri et al., 2019, 2022a). A detailed discussion of various viscoelastic models can be found in Jackson & Faul (2010); Faul & Jackson (2015); Bagheri et al. (2019). We consider both Maxwell and Andrade viscoelastic models because of the potential proximity of the relaxation time of the ice shell to Enceladus’s orbital period, but most of our results assume the Andrade rheology.
2.3 Physical libration
Owing to Enceladus’s moderate orbital eccentricity and slightly elongated shape, it is subject to periodic torques that force harmonic oscillations in its orientation, called physical librations. This libration is superimposed on the overall synchronous orbit of the moon around Saturn. The amplitude of physical libration depends mainly on the moments of inertia of the body, and whether the outer surface layer–the ice shell–is coupled to the deep interior (e.g., Rambaux et al., 2010), and to less extent to the rigidity of the shell. The estimated libration of Enceladus based on the analysis of observations made by Cassini is too large to be explained by an ice shell that is mechanically coupled to the core (Thomas et al., 2016). Recent work by Park et al. (2024) suggests a slightly lower amplitude of libration, although still higher than expected if the ice shell and the core are coupled. A global ocean which mechanically decouples the shell from the rocky core is inferred from the libration amplitude (Thomas et al., 2016). It should be noted that the shell and core might decouple without a global ocean in the presence of another frictionless boundary layer, such as found in soft-bedded glacier systems on Earth (Johnston & Montési, 2017). While the scenario of a non-global ocean cannot be ruled out by the constraint from libration, here we assume the presence of a global subsurface ocean.
Tides can influence the amplitude of libration by various mechanisms: by modifying the gravitational torque exerted by the planet on the satellite; through gravitational coupling of the libration of the ice shell to the solid interior through the torque between the periodic tidal bulge of one solid layer and the permanent shape of another; and by the slight periodic modification of the polar moment of inertia by the zonal tides.The amplitude of libration of the outer ice shell is given by (Van Hoolst et al., 2013):
$\gamma_{I}=\frac{1}{C_{c}C_{s}}\frac{4e[K_{3}K_{5}+K_{2}K_{6}-n^{2}C_{c}K_{3})]}{(n^{2}-\sigma_{1}^{2})(n^{2}-\sigma_{2}^{2})},$ (7)
where $K_{1}$, $K_{2}$, $K_{3}$, $K_{4}$, $K_{5}$ and $K_{6}$ describe the torques between the planet, the shell, and the core, $C_{c}$ and $C_{s}$ are the polar moments of inertia of the core and the shell, $\gamma_{c}$ and $\gamma_{I}$ are the amplitudes of libration of the core and the shell, and $\mathcal{M}$ is the mean anomaly, and
$\sigma_{1,2}^{2}=\frac{K_{1}C_{c}+K_{5}C_{s}\pm\left[4(K_{2}K_{4}-K_{1}K_{5})C_{c}C_{s}+(K_{1}C_{c}+K_{5}C_{s})^{2}\right]^{1/2}}{2C_{c}C_{s}}.$ (8)
Equation (8) represents the natural frequencies of free libration of the body (Van Hoolst et al., 2013). In order to compute the flattening of each of the layers, as required for the calculation of the torques between layers in equations (78), we use the approach of Tricarico (2014) that provides hydrostatic equilibrium figures for moons in eccentric orbits, along with the global shape from the best-fitting ellipsoid model suggested by Park et al. (2024).
2.4 Static gravity field
Current gravity data for Enceladus is limited to that obtained from three Cassini fly-bys. Therefore, we are limited to gravity field coefficients of spherical harmonic degree $l=2,3$. To further constrain the interior structure, we use the moment of inertia and the $l=3$ zonal potential coefficients ($J_{3}$). We do not use the $l=2$ gravity coefficients ($J_{2}$ and $C_{22}$) because of the underlying assumptions required to account for the non-hydrostatic component of the gravity field and the compensation mechanism, as discussed in Hemingway & Mittal (2019). Direct measurement of the moment of inertia of Enceladus is possible only by measuring the precession of the body. Hemingway et al. (2018) propose to derive the moment of inertia from long-wavelength $l=2$ gravity-topography admittance measurements ($Z_{22}$ and $Z_{20}$). We refer the reader to McKinnon (2015), Hemingway et al. (2018), and Hemingway & Mittal (2019) for the details of this derivation. Here, we only consider the moment of inertia when we explore the potential impact of future observations. To compute the zonal non-hydrostatic degree-3 static gravity coefficient ($J_{3}$), we follow an approach similar to that described in Genova et al. (2024) to compute the zonal harmonic gravity coefficient. The gravity-topography admittance is computed as (Turcotte et al., 1981):
$Z_{l}=4\pi G\bigg{(}\frac{l+1}{2l+1}\bigg{)}\rho_{\rm shell}\Bigg{[}1-\bigg{(}1-\frac{D_{shell}}{R_{E}}\bigg{)}^{l+2}\bigg{(}\frac{g_{t}}{g_{b}}\bigg{)}C_{l}^{0}\Bigg{]}.$ (9)
Here, $D_{shell}$ is the mean shell thickness, $C_{l}^{0}$ is a compensation factor computed based on the effective elastic thickness $D_{e}$ (Turcotte et al., 1981), and $g_{t}$ and $g_{b}$ denote the gravitational acceleration at the top and bottom of the shell:
$g_{t}=\frac{GM}{R_{E}^{2}},\$
$g_{b}=\frac{GM}{(R_{E}-D_{\rm shell})^{2}}-\frac{4}{3}\pi G\rho_{\rm shell}\frac{R_{E}^{3}-(R_{E}-D_{\rm shell})^{3}}{(R_{E}-D_{\rm shell})^{2}}.$ (10)
$J_{3}$ is computed as:
$J_{3}=\frac{Z_{3}R_{E}^{2}H_{30}^{n\mathrm{hyd}}}{4GM}.$ (11)
$H_{30}^{nhyd}$ is the non-hydrostatic topography coefficient. This method implies an unconstrained degree of compensation. Following Genova et al. (2024), we assume Airy compensation, implying $C_{l}^{0}=1$. This assumption does not affect the uncertainties we derive for the observations and model parameters.
3 End-member scenarios for tidal dissipation in Enceladus
The viscosity of consolidated rock is higher than $10^{18}$ Pa.s (e.g., Hirth & Kohlstedt, 2003). However, unconsolidated material, or even a mixture of rock and ice, may have a significantly lower effective viscosity. In this spirit, Choblet et al. (2017) propose a range of lower core viscosities that allows substantial tidal heating in the core of Enceladus. Alternatively, tidal dissipation in the icy crust may explain the tidal heating at Enceladus if the temperature of ice is sufficiently warm, and thus the effective shell viscosity is sufficiently low. The estimated tidal heating from these different scenarios is shown in Figure 1(a). The four narrow regions determined by the red lines indicate the ranges of the estimated tidal heating in Enceladus and the associated effective viscosities of the shell and the core, assuming that the tidal heating is fully balanced by the heat loss rate at the surface in the range of 15–40 GW (Howett et al., 2011; Hemingway & Mittal, 2019; Nimmo et al., 2023; Park et al., 2024). We show the variation in tidal heating with the viscosity of the shell (core) in Figure 1 panels (b) and (c) for a fixed value of core (shell) viscosity. Dissipation peaks at certain viscosities of both the shell and core at which the associated Maxwell relaxation time is close to the period of excitations, and the tidal response transitions from anelasticity-dominated behavior to viscous-dominated behavior. Therefore, the estimated tidal dissipation rate is compatible with two regions around the peak dissipation (highlighted in gray). For the interior structure assumed here, we find that an effective core viscosity O(1013) Pa.s can satisfy the total estimated heat, consistent with that suggested by Choblet et al. (2017).
The two end-member scenarios manifest themselves differently in the absolute values of $k_{2}$, $h_{2}$, and $l_{2}$ and their respective phase lags (Figure 2). When the effective viscosity of the rocky core, $\eta_{core}$, is $\lesssim 10^{13}$ Pa.s., deformation of the core will significantly affect the amplitude of  $k_{2}$ and its phase lag $\epsilon_{k_{2}}$ (Figure 2 panels (a) and (d)). However, tidal deformation of the core does not substantially affect tidal displacement at the surface, resulting in only a small change in $h_{2}$. This small influence on $h_{2}$ is found to be less than 10% for Enceladus. The compatibility of the steady-state heat budget with both dissipation scenarios presents a degeneracy that requires additional constraints to resolve. In the case of Enceladus and for most of the icy bodies where the ice shell is mechanically decoupled from the solid interior by an ocean, the ambiguity in the location of the tidal heating can be resolved by considering the tidal response at the surface and the associated phase lag. In the case of a mechanically coupled interior or tidally inactive interior layers, there is a strong correlation between the amplitude and phase lag of deformations ($h_{2}$, $l_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{l_{2}}$) on the surface and those of the observed gravitational signal. However, a lack of such a correlation would imply that tidal activity in deeper interior layers decoupled from the surface, such as the core, is causing a perturbation in the gravitational field of the body. In other words, if the ice shell is the primary locus of tidal heat generation, then the shell undergoes considerably larger deformations with a significant phase lag that incorporates viscoelastic deformations, as opposed to the case where the tidal activity occurs predominantly in the rocky core and the deformations in the shell remain largely elastic. For a given interior structure model of Enceladus, the tidal deformations on the surface in the two cases differ by approximately a factor of two (Figure 2 panels (b) and (c)). The phase lags associated with tidal displacements can differ by tens of degrees between the two cases.
Ermakov et al. (2021) and Genova et al. (2024) propose suites of geophysical measurements to study the interior of Enceladus. However, the ranges of elastic and viscoelastic parameters they consider for the rocky core is such that tidal deformation and dissipation in the core are effectively negligible. Ermakov et al. (2021) assume $\eta_{core}>10^{20}$ Pa.s, $\mu_{core}>10$ GPa , and Genova et al. (2024) assume $\eta_{core}>10^{16}$ Pa.s, $\mu_{core}>50$ GPa. The results of these studies represent only one of the tidal dissipation scenarios, i.e., tidal dissipation occuring in the shell without effective contribution from the core. Here, we broaden the parameter space to allow for the possibility of significant tidal dissipation in the core. Note that the commonly assumed viscosity of ice, as suggested by Goldsby & Kohlstedt (2001) ($\eta_{\rm ice}\approx 10^{14}$ Pa.s,) corresponds to ice somewhat near its melting point. Due to the strong temperature dependence of viscosity and very low temperature conditions at Enceladus that may persist deep within the shell, assuming a uniform viscosity of $10^{14}$ Pa.s for the shell (Efroimsky, 2018; Ermakov et al., 2021; Genova et al., 2024) is unrealistic.
Figure 1: (a) Tidal heating (in GW) versus the effective viscosity of the ice shell and the core. The narrow regions bounded by the red lines indicate the range of heat loss that covers the estimates by Howett et al. (2011); Hemingway & Mittal (2019); Park et al. (2024). (b) Tidal heating variation with ice shell effective viscosity for a fixed core viscosity of 1020 Pa.s. (c) Tidal heating variation with core effective viscosity for a fixed shell viscosity of 1016 Pa.s. The gray areas highlight the region where both viscosity and the tidal heating values are compatible with the observed steady state heating. The results are presented for $D_{\rm shell}=14~{}$km, $D_{\rm ocean}=34$ km, $\rho_{\rm ocean}=1000~{}\rm kg/m^{3}$, $R_{\rm core}=204$ km, $\rho_{\rm shell}=930~{}\rm~{}kg/m^{3}$, $\mu_{\rm shell}=3$ GPa, $\kappa_{\rm shell}=10$ GPa, $\rho_{\rm core}=2175~{}\rm~{}kg/m^{3}$, $\kappa_{\rm core}=20$ GPa, $\mu_{\rm core}=15$ GPa. We have assumed an Andrade viscoelastic model.
Figure 2: Tidal Love number amplitudes (panels a, b, c), phase lags (d, e, f) for different shell and core viscosities. Same interior structure parameters as in Figure 1 are used.
As described in Section 2.1, there exist various viscoelastic models based on theoretical and laboratory results. Since the nominal values of the effective relaxation time of ice could be close to the orbital period (tidal forcing period) of Enceladus, we examine the choice of Maxwell and Andrade viscoelastic models by computing the tidal response components (Love numbers and heat) using both models. Our analysis shows that if the viscosity of the ice shell is sufficiently high due to the low temperature ($\eta_{\rm shell}\gtrsim 10^{18}$ Pa.s), its relaxation time will be significantly longer than the forcing period, implying a predominantly elastic behavior of the shell, and therefore the choice of the viscoelastic model does not significantly affect the response (Figure 16 in Appendix A). However, if the viscosity of the ice shell is such that the relaxation time ($\tau_{M}~{}=~{}\eta_{\rm shell}/\mu_{\rm shell}$) is close to the forcing period, that is, $\eta_{\rm shell}~{}\approx~{}10^{13-14}$ Pa.s, the Maxwell and Andrade viscoelastic models predict very different surface deformation ($\sim$3 m, see Figure 16 in Appendix A). The two viscoelastic models differ more significantly in their predicted imaginary component of the response, and thus the generated frictional heat, as opposed to the real part of the response which is largely determined by the rigidity of the body. The difference in total tidal heating generated in both the ice shell and the core, following the difference in the $k_{2}$ amplitude and the phase, can reach several GW. Although we do not rule out the applicability of the Maxwell viscoelastic model (see e.g., Hussmann et al., 2016), for further calculations in this paper, we adopt the Andrade model because of its capability to consider transitional anelastic mechanisms missing in the Maxwell model. The extent to which Enceladus’s tidal response deviates from a Maxwell body to an Andrade or more complex viscoelastic body is determined by the large-scale rheological state of its interior (see e.g., Bagheri et al. (2022b); Bierson (2024) for discussions on various viscoelastic models). We investigate the sensitivity of Enceladus’s tidal response to the choice of Andrade parameters in Appendix A.
4 Sensitivity of the ice shell thickness to physical libration and tidal coupling
Using the formulation provided in Section 2.3 and the same physical parameters used in Figures 1 and 2, we consider the sensitivity of the amplitude of libration to the thickness of the ice shell and the effective viscosity of the shell and the core (Figure 3). We use a range for the amplitude of the libration of 360 m$<\gamma_{I}<$750 m, which spans the estimates of Nadezhdina et al. (2016), Thomas et al. (2016), and Park et al. (2024). The sensitivity of the libration to the shell thickness is not uniform for different ice shell thickness ranges. Much higher variations in the libration amplitude are observed when the mean thickness of the ice shell is close to 5–10 km. At these thicknesses, the natural frequency of Enceladus approaches the frequency of the physical libration, resulting in a resonance that significantly enhances the libration amplitude. All current estimates of the physical libration of Enceladus are lower than those suggesting nearly resonant behavior, which rules out a mean ice shell of thickness 5–10 km. The libration amplitude is less sensitive to the viscosity of the ice shell compared to its thickness. However, for ice shell viscosities near the nominal relaxation viscosity of the ice, the effect of the shell’s viscoelastic deformation should not be ignored. This effect is approximately 10% if the viscosity of the shell is as low as $\sim$12.5$\times$1013 Pa.s compared to high-viscosity ice ($\gtrsim 10^{15}$ Pa.s) where there is negligible impact of anelasticity on the predicted libration. As expected, libration is much less sensitive to the viscosity of the core because the only effect from the core is through the gravitational perturbations caused by the core’s tidal deformation on the ice shell, which are insignificant even in the case of a tidally active core.
Figure 3: Variation of the amplitude of the libration as a function of ice shell thickness and the effective viscosities of shell (a) and the core (b). In panel (a), the core viscosity is assumed equal to $\rm 10^{20}~{}Pa.s$. The region between the red lines shows the libration amplitude spanning the estimates by Thomas et al. (2016); Nadezhdina et al. (2016); Park et al. (2024), i.e., ($\rm 360~{}m<\gamma_{I}<750~{}m)$. Other interior structure parameters are the same as in Figure 2. At $\rm D_{~{}shell}\approx 5$ km, the natural free libration period of the shell approaches the orbital period of Enceladus. The x-axes and the libration color scale are logarithmic.
Quantity Symbol Unit Value
Radius $R_{E}$ km     251.99${}^{(1)}$
Mass (×$\times 10^{11}$) $M$ kg $1.08101\pm 1.08~{}^{(2)}$
Flattening coeff.(×$\times 10^{6}$) $C_{20}$   –5564.4 $\pm$ 37.58${}^{(1)}$
$-5526.1\pm 35.5~{}^{(3)}$
Quadruple coeff.(×$\times 10^{6}$) $C_{22}$     1542.0 $\pm$ ${}^{(1)}$
$1575.7\pm 15.9~{}^{(3)}$
Degree-3 coeff.(×$\times 10^{6}$) $J_{3}$ $-180.64\pm 33.95~{}^{(1)}$
  –118.2 ±$\pm 23.5~{}^{(3)}$
SPT surface heat flux $\dot{E}_{SPT}$ GW       $15.8\pm 3.1~{}^{(4)}$
Estimated total surface heat flux $\dot{E}_{total}$ GW        $18-28~{}^{(1)}$
          $25-40~{}^{(5,6)}$
Physical libration $\gamma_{I}$ deg         $0.091\pm 0.009~{}^{(1)}$
          $0.12\pm 0.014~{}^{(7)}$
        $0.155\pm 0.014~{}^{(8)}$
Orbital eccentricity $e$        $0.0047~{}^{(9)}$
Orbital mean motion (×$\times 10^{-5}$) $\omega$ rad/s        5.307${}^{(10)}$
Table 1: Existing measurements for Enceladus. The subscripts indicate the references for the quoted values: 1Park et al. (2024), 2Jacobson (2022), 3Hemingway et al. (2018), 4Howett et al. (2011), 5Nimmo et al. (2023), 6Hemingway & Mittal (2019), 7Thomas et al. (2016), 8Nadezhdina et al. (2016), 9Nimmo et al. (2018), 10Roberts & Nimmo (2008). Static gravity coefficients have been adjusted to a reference radius of 251.99 km with respect to the calculated values in Park et al. (2024) that used a radius of 256.6 km.
Quantity Symbol Unit Range
Shell thickness $D_{\rm shell}$ km $2-50$
Shell effective viscosity $\eta_{\rm shell}$ Pa.s $10^{10}-10^{25}$
Shell shear modulus $\mu_{\rm shell}$ GPa $2-4$
Shell bulk modulus $\kappa_{\rm shell}$ GPa 10 (fixed)
Ocean density $\rho_{\rm ocean}$ kg/m3 $930-1400$
Ocean thickness $D_{\rm ocean}$ km $1-130$
Ocean bulk modulus $\kappa_{\rm ocean}$ GPa 2.1
Core radius $R_{\rm core}$ km $120-220$
Core shear modulus $\mu_{\rm core}$ GPa $1-50$
Core bulk modulus $\kappa_{\rm core}$ GPa 100 (fixed)
Core density $\rho_{\rm core}$ kg/m3 $2000-2800$
Core viscosity $\eta_{\rm core}$ Pa.s $10^{9}-10^{22}$
Table 2: Prior ranges for interior structure model parameters. We sample all the model parameters on logarithmic scale.
5 Constraining the partitioning of tidal heating between the shell and the core
The total tidal energy dissipated in Enceladus can be computed by measuring the amplitude and phase lag of $k_{2}$, independent of the choice of rheological model (see Equation (5)). We can compute the tidal heating in any specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body. Here, we use equation 37 in Tobie et al. (2005) to derive the contribution of the core and the shell to the global tidal heating:
$H_{tidal}(r)=-\frac{21}{10}\frac{\omega^{5}e^{2}R_{E}^{4}}{r^{2}}H_{\mu}\Im{(\mu(r))},$ (12)
where $H_{tidal}$ is the radial distribution of the dissipation rate per unit volume, $r$ is the radius from the center, $H_{\mu}$ is the sensitivity kernel for shear dissipation given in Equation (33) in Tobie et al. (2005), and $\Im{(\mu(r))}$ is the imaginary part of the complex shear modulus which is assumed constant at each layer. By integrating this expression from the center of the satellite to the core-ocean boundary, we obtain:
[…]
The integral can be written in terms of the imaginary part of the piecewise tidal Love number of the core following Equation 35 in Tobie et al. (2005):
[…]
Here, $\hat{k}_{2}^{core}$ is the piecewise gravitational Love number associated with the core. Thus, if we compute the contribution of each of the layers to $|k_{2}|/Q$ or identically to $\Im(k_{2})$ of the body, that is, evaluating at each interface, we can compute the associated tidal heating for the part of the body that is interior to the interface. For a homogeneous core, $\hat{k}_{2}^{core}$ can be computed from the deformation at the top of the core:
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}}$ (15)
Tidal heating in the ice shell is then computed by subtracting the contribution of the core from the total energy dissipation:
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (16)
Equations (13) and (16) provide a method to calculate the partition of the global tidal heating for a given interior structure model. These equations require a priori assumptions or independent constraints on the rheological and structural properties to compute the shear dissipation in the core. Therefore, here we derive the partitioning of tidal heating on the basis of the observable quantities, i.e., the tidal Love numbers. We extend the formulation provided by Beuthe (2019) which uses membrane mechanics to obtain the contribution of each layer of the body to the imaginary part of the tidal Love number. The shell’s contribution can be written as (Beuthe, 2019):
$\Im{(\hat{k}_{2}^{shell})}=\left|\frac{k_{2}+1}{k_{2}^{0}+1}\right|^{2}\Im{(k_{2}^{0})}-\xi_{2}|h_{2}|^{2}\Im{(\Lambda_{T})},$ (17)
where, following the terminology used in Beuthe (2019), $\xi_{2}=3\rho_{ocean}/5\rho_{E}$, $\rho_{E}$ is the density of the whole body, $k_{2}$ and $h_{2}$ are the tidal Love numbers of the whole body, and $k_{2}^{0}$ is the fluid-crust gravitational Love number (assuming the crust is fluid), and $\Lambda_{T}$ is the shell spring constant introduced by Beuthe (2019). In Appendix B, we show that the tidal dissipation occurring in the shell can be expressed in terms of the tidal Love numbers and density of the layers:
$\dot{E}_{\text{shell}}=\frac{105}{10}\frac{(\omega R)^{5}}{G}\xi_{2}\,|h_{2}|^{2}\bigg{[}\Im\bigg{(}{\frac{1}{h_{2}}\big{(}k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon\big{)}}\bigg{)}\bigg{]}\,e^{2}.$ (18)
Here, $\delta\rho$ is the density jump from the ocean to the crust and $\epsilon$ is the ratio of the thickness of the crust to Enceladus’s radius. Equation (18) expresses the tidal heating in the shell using measurements of the real and imaginary parts of the Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers as required in Equation (17). Tidal heating in the core can be accordingly computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
6 Road-map to Constrain the interior properties with geodetic observations
Despite numerous studies based on the observations made by Cassini, multiple interior structure properties of Enceladus are not well constrained. This uncertainty is due in part to the limitation of the available observations and in part to differences in the interpretation of the data for the interior structure and the underlying modeling assumptions (e.g., Howett et al., 2011; Spencer et al., 2013; Thomas et al., 2016; Park et al., 2024; Schenk & McKinnon, 2024). Here, we explore the ranges of possible interior parameters that are consistent with the currently available observations. Based on these constraints, we predict ranges of the currently unavailable geodetic observables, and investigate the possibility of discriminating between the two main tidal heating scenarios from future measurements. We use the estimates on the observables as a basis for exploring the interior from future measurements.
6.1 Constraining the interior and discriminating between heat generation scenarios
Currently available observations that can be used to understand Enceladus’s interior structure and heat production include shape (e.g., Park et al., 2024; Schenk & McKinnon, 2024), gravity up to degree-3 (Iess et al., 2014; Park et al., 2024), physical libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024), and heat flux (Howett et al., 2011; Spencer et al., 2013) as summarized in Table 1.
Figure 4: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and libration amplitude spanned by estimates in Thomas et al. (2016), Nadezhdina et al. (2016) and Park et al. (2024). Other observables in this plot are free to vary. In the 2D Probability Density Functions (PDFs), black (yellow) color indicates low (high) probability. The complex patterns observed indicating multiple peaks in the PDFs of $k_{2}$, $h_{2}$, and $l_{2}$ and phase lags reflect the different regimes of dissipation consistent with the observations (see Figure 2), resulting in the nonlinearity of the forwards model. Phase lags of $k_{2}$, $h_{2}$, and $l_{2}$ are in degrees and amplitude of the libration is in meters. $J_{3}$ is shown divided by 106.
We start with the most conservatively chosen set of observational constraints to show the effect of incorporating various sets of observations on the measurement requirements for future exploration. We assume spherical symmetry and consider a set of interior parameters, i.e., thickness, rigidity, effective viscosity, and density of the layers, and geodetic observable quantities, i.e., tidal Love numbers, phase lag angles, libration, moment of inertia, and $l=3$ zonal gravity coefficient, hereafter referred to as observables. We conduct Monte Carlo sampling based on the prior ranges of the model parameters presented in Table 2 (see Appendix C). We present the correlation between the observables in Figure 4. We use the total density, the physical libration amplitude in a range covering the estimates by Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024), and the estimated heat loss rate as the only constraints. To demonstrate the widest potentially plausible parameter space, we have also considered a wide range of published estimates of the heat loss rate suggested in Howett et al. (2011), Hemingway & Mittal (2019), and Park et al. (2024) (15 GW$<\dot{E}<$40 GW). The Love numbers, phase lags, moment of inertia, and degree-3 static gravity coefficient vary freely as long as the model satisfies the constraints of mass, heat loss, and libration; that is, we do not impose any constraints on the estimated moment of inertia and $J_{3}$ to avoid any possible inconsistency of our models with the underlying assumptions used in deriving these estimates.
Consistent with Section 3, the patterns of the posterior Probability Density Functions (PDFs) of the amplitude and phase lag of $k_{2}$ largely differ from the PDFs of the amplitudes and phases of $h_{2}$ and $l_{2}$ (Figure 4). PDFs of $k_{2}$ and $\epsilon_{k_{2}}$ have several peaks that are associated with tidal activity dominated in the core versus in the shell. A different pattern of distribution and peaks associated with different regimes of tidal activity is observed for $h_{2}$ and $l_{2}$, and their phase lags, since these quantities are largely insensitive to tidal activity in the core. The posterior PDFs show that while the $h_{2}$ and $l_{2}$ and their phase lags (Figure 4) are strongly correlated with each other (dominated by the response of the shell), their correlation with the amplitude and phase lag of $k_{2}$ depends on where the tidal dissipation dominantly occurs.
We also obtain the ranges of the interior model parameters and the correlation between them associated with the observables (Appendix Figure 18). Previously, we ruled out mean crustal thicknesses around 5–10 km for Enceladus, as they would result in amplitudes of libration that are too large due to resonance in the shell (see Section 4). Ice shell thicknesses $\approx$ 1–5 km would be compatible with the estimates of tidal heating and libration only if the mean effective viscosity of the shell is unrealistically low. Therefore, we conclude that the mean thickness of Enceladus’s shell is larger than 10 km. For completeness, we included very low effective viscosities for the shell and the core, which imply deformations in the viscous-dominated regimes. As a result, the PDFs of the viscosities of the shell and core indicate multiple peaks representative of tidal dissipation dominated by viscous deformation as well as the anelastic deformation (See also Figure 1). From here on out, we consider viscosities of the ice shell and the rocky core that are close to or greater than their nominal relaxation viscosities ($\eta_{\rm shell}~{}>~{}10^{12}$ Pa.s and $\eta_{\rm core}~{}>~{}10^{11}$ Pa.s), i.e., the ice shell and the rocky core deform in the anelasticity-dominated deformation regime and not in the viscous-dominated deformation regime.
We focus on analysis of discriminating between two sets of geodetic observables and interior parameters that are associated with each of the two scenarios (Figures 5, and Figures 6). We compute the tidal heating in the shell using Equation (16). We differentiate between the core-dominated and shell-dominated tidal dissipation scenarios by defining a threshold whereby at least 90% of the total heat is produced in the shell or the core. The PDFs of $k_{2}$ and $\epsilon_{k_{2}}$ for the two tidal dissipation scenarios largely overlap (Figures 5 and 6), whereas the PDFs of $h_{2}$, $\epsilon_{h_{2}}$, $l_{2}$, and $\epsilon_{l_{2}}$ in the case of shell-dominated tidal dissipation (Figure 5) differ significantly from those in the case of core-dominated tidal dissipation (Figure 6). In an inverse problem context, the degeneracy caused by the two different tidal dissipation scenarios reflected in the tidal Love numbers, as discussed earlier, leads to the non-uniqueness of the plausible interpretation of $k_{2}$ and $\epsilon_{k_{2}}$.
We obtain tight constraints on the viscosity of the shell in the case where we impose the constraint that the tidal activity is dominated in the shell, and on the viscosity of the core in the case where a core-dominated tidal dissipation occurs (Figure 7). The rheological properties of the region in which tidal dissipation is not focused remain largely unconstrained. Using this set of observables, we find that the thickness of the ice shell, mostly constrained by the libration amplitude, is in the range of 15–40 km. Since we have conservatively chosen broad ranges for observables in this analysis, we have not obtained tight constraints on most of the other model parameters in this analysis. The PDFs and correlations between the interior structure parameters are presented in Appendix Figures 19 and 20.
Figure 5: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and the libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024). Here, we assume that the dissipation is dominantly occurring in the ice shell. Unit conventions are the same as in Figure 4.
Figure 6: Exploration of the potentially observable geodetic quantities consistent with mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024). Here, we assume that tidal dissipation occurs dominantly in the core. Unit conventions are the same as in Figure 4.
Figure 7: PDFs of ice shell and core viscosities for each of the two cases of tidal dissipation occurring dominantly in the shell (purple) and in the core (yellow), corresponding to Figure 5 and Figure 6, respectively.
We compare the PDFs of $h_{2}$ and $l_{2}$ and their phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) associated with the two heat generation scenarios in Figure 8. The 2-$\sigma$ standard deviations are indicated in this figure. The separation between the two intervals associated with the 2-$\sigma$ uncertainties corresponding to the two PDFs determines the measurement precision required to discriminate between the two dissipation regimes. In other words, if we measure $h_{2}$ with a precision equal to the separation between the 2-$\sigma$ confidence intervals (or 1-$\sigma$, not shown in the figure) of the two PDFs, we obtain a confidence $\sim$95(or 68% if 1-$\sigma$ intervals are used) in identifying the layer that dominates the dissipation budget. Due to the overlap between the PDFs of $h_{2}$ and $l_{2}$ associated with the two tidal dissipation scenarios using this set of observations, distinguishing between the two cases is challenging. This large overlap is due to the large uncertainties in the observables considered in this case, and taking into account only the currently available constraints. The analysis implies that measurement of the maximum vertical tidal deformations with a maximum uncertainty of 62 cm can distinguish between the dominant region of tidal activity within the 1-$\sigma$ confidence level. Measurement of radial deformations with a maximum uncertainty of 56 cm can indicate the more “likely” region of heat generation, but it cannot definitively discriminate between the two scenarios (This value is obtained by the separation associated with the 2-$\sigma$ uncertainties between the two regimes which overlap in this case). Alternatively, measurement of the tangential displacements within an uncertainty of 28 cm is required for 1-$\sigma$ confidence. The phase lags associated with the radial deformations ($\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$) indicate a better separation between the PDFs associated with the two regimes of tidal activity compared to the deformations and $h_{2}$. We find that measuring the phase lag angles of the tidal deformations with an accuracy of 11 is required for 1-$\sigma$ confidence, and 4 is required for 2-$\sigma$ confidence in determining the dominant region of tidal dissipation.
Figure 8: Vertical and horizontal displacement Love numbers ($|h_{2}|$, $|l_{2}|$) and phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) are presented for the core-dominated and shell-dominated tidal dissipation. The heat flux is assumed to be $\rm 15~{}GW<\dot{E}<40~{}GW$, which includes the lowest and highest ranges estimated in Hemingway et al. (2018) and Park et al. (2024), and the direct measurements (Howett et al., 2011). Libration is conservatively assumed to be between $\rm 0.082^{\circ}<\gamma_{I}<0.169^{\circ}$, covering the estimates in Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024). The top axis in the plots for $h_{2}$ and $l_{2}$ indicate the associated maximum radial and maximum easterly displacements corresponding to the values of the Love number in the bottom axis, respectively.
Next, we focus on a narrower range of observational constraints using exclusively the values and uncertainties in Park et al. (2024) for $\gamma_{I}$, $J_{3}$, moment of inertia, mean density, and the total heat loss rate (see Table 1) to constrain the interior structure. We use 3-$\sigma$ uncertainties on the observables. We present the correlation between the observables and the model parameters obtained based on this analysis for a case that incorporates both tidal dissipation scenarios in Figures 9a and 9b. The obtained preferred values of the observables and uncertainties based on Park et al. (2024), including the tidal Love numbers, phase lags, moment of inertia, libration amplitude, and $J_{3}$ are given in Table 3. We also quantify the values for the observables that present the highest likelihood for each of the two heat generation scenarios. We will use these synthetic observations to perform inversions using plausible uncertainties expected from future measurements. Using the narrower observational constraints based on Park et al. (2024) further demonstrate that while $k_{2}$ and $\epsilon_{k_{2}}$ show a normal distribution with a single peak, $h_{2}$, $l_{2}$, $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ show two peaks associated with the two tidal dissipation scenarios. We note that satisfying all the measurements provided in Park et al. (2024) is possible when using 3-$\sigma$ uncertainties on the observables. Specifically, we find that the mean value of $J_{3}$ can be matched with the libration amplitude when considering 3-$\sigma$ uncertainty on the measured $J_{3}$. In contrast, Genova et al. (2024) use only the gravity measurements without exploiting the libration measurement as a constraint, which is expected to somewhat alter the results. Our results can also be compared to those presented by Genova et al. (2024) and Ermakov et al. (2021) where only tidal heating in the shell is considered. We also obtain considerably improved constraints on the interior parameters (Figure 9b) compared to the case where we use broader uncertainties on the observations. However, several model parameters remain weakly constrained with the currently available observations. The ranges of the obtained plausible interior parameters are listed in Table 4. See Appendix Figures 21 and 22 for the 2D PDFs of the model parameters associated with each of the cases.
Observable Unit Dissipative core Dissipative shell Expected Uncertainty
$|k_{2}|$ 0.0240 $\pm$ 0.0102 0.0317 $\pm$ 0.0130 0.002 (0.0002)
$|h_{2}|$ 0.0459 $\pm$ 0.0294 0.0848 $\pm$ 0.0359 0.0172
$|l_{2}|$ 0.0104 $\pm$ 0.0072 $0$$\pm$ 0.0090
$\epsilon_{k_{2}}$ deg 39.35 $\pm$ 19.72 28.91 $\pm$ 17.25 4 (0.4)
$\epsilon_{h_{2}}$ deg 2.95 $\pm$ 2.69 28.73 $\pm$ 17.18
$\epsilon_{l_{2}}$ deg 3.10 $\pm$ 2.85 29.71 $\pm$ 17.89
$J_{3}(\times 10^{-6})$ 260.4 $\pm$ 36.9 259.3 $\pm$ 38.06 0.4 (0.04)
$\gamma_{I}$ m 413.4$\pm$ 65.9 408.1 $\pm$ 68.6 4.0 (1.3)
Total heat GW 18–28 18–28
Density kg/m3 1611 $\pm$ 1.6 1611 $\pm$ 1.6
Table 3: Favored values and uncertainties (3-$\sigma$) of observables calculated using MCMC sampling based on the available inferred observables given by Park et al. (2024). Also, expected uncertainties from future explorations are provided (see Section 6.4) which are used in the inversion in Section 6.2. The values in the parentheses in the expected measurements of the tidal Love numbers and phase lag corresponds to a scenario with lower accuracy, as discussed in the text.
Quantity Unit Tidally active core Tidally active shell
$D_{\rm shell}$ km $22-34$ $22-34$
$\eta_{\rm shell}$ Pa.s $6.3\times 10^{14}-10^{20}$ $2.5\times 10^{13}-1.3\times 10^{14}$
$\mu_{\rm shell}$ GPa $2-4$ $2-4$
$\rho_{\rm shell}$ kg/m3 $700-950$ $700-950$
$\rho_{\rm ocean}$ kg/m3 $920-1250$ $920-1250$
$R_{\rm core}$ km $183-213$ $183-213$
$\mu_{\rm core}$ GPa $4-70$ $5-80$
$\rho_{\rm core}$ kg/m3 $2160-2400$ $2160-2400$
$\eta_{\rm core}$ Pa.s $4.4\times 10^{11}$$2.3\times 10^{12}$ $3.16\times 10^{13}-10^{22}$
Table 4: Ranges of the model parameters from two of MCMC analysis satisfying the currently available measurements based on the analysis of observables in Park et al. (2024).
The tighter constraints obtained from Park et al. (2024) allow relaxation of the demands on measuring $h_{2}$ and $l_{2}$ in order to discriminate between the two heating scenarios. Based on the observational constraints of Park et al. (2024) (as in Figure 9a), we show the amplitudes and phase lags of the Love numbers in Figure 10 panels (a)-(f). Using this narrower set of observational constraints, the PDFs of the tidal response associated with the two tidal dissipation scenarios are further separated compared to Figure 8 where broader ranges of values for the observables were assumed. In this case, measuring the maximum vertical deformations with a precision of 87 cm is sufficient to differentiate between the two hypotheses of heat generation with a 1-$\sigma$ confidence. For a confidence level of 2-$\sigma$, measuring the maximum radial displacement of 1 cm is required. Alternatively, measuring the phase lag $\epsilon_{h_{2}}$ by 20 and 15 is sufficient to discriminate between the two tidal dissipation scenarios with 1-$\sigma$ and 2-$\sigma$ confidences, respectively. Also, our analysis indicates that measurements of maximum tangential displacements within 40 cm and 0.5 cm are required to differentiate between the two tidal dissipation scenarios providing 1-$\sigma$ and 2-$\sigma$ level confidences, respectively. Similarly to $\epsilon_{h_{2}}$, the measurement precisions required for $\epsilon_{l_{2}}$ for the confidence levels of 1-$\sigma$ and 2-$\sigma$ are 20 and 15, respectively.
In case future measurements provide constraints on $k_{2}$, the measurement requirement for $h_{2}$ and/or $l_{2}$ could be significantly relaxed. Therefore, we present PDFs of displacement tidal Love numbers associated with the core- and shell-dominated tidal dissipation assuming additional constraints on the amplitude and phase lag of $k_{2}$ (Figure 10 panels (g)-(j)). If the real and imaginary parts of $k_{2}$ are measured within 0.002, the accuracies required to measure the maximum radial deformations to obtain 1-$\sigma$ and 2-$\sigma$ confidence levels relax to 134 cm and 82 cm, respectively. In that case, the upper bounds of the uncertainty required for the horizontal deformations are 61 cm and 36 cm, respectively. The accuracies required for the phase lags of $h_{2}$ or $l_{2}$ are both equal to 21 and 16, for confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively.
Furthermore, in case a more precise measurement of $k_{2}$ (both amplitude and phase) is obtained, i.e. within an uncertainty of 0.0002, the required measurement of radial or horizontal deformations will be further relaxed (not shown here). In that case, the requirement for the maximum radial displacement will be 150 cm and 116 cm for the 1-$\sigma$ and 2-$\sigma$ confidence levels, respectively. These values are 68 cm and 51 cm for the tangential component of the deformation, respectively. The precisions required for the phase lags of $h_{2}$ or $l_{2}$ are equal to 24 and 22 to obtain confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively. We summarize the precision required for measuring the radial and horizontal displacements based on different scenarios discussed here in Table 5. Note that the measurements of the $h_{2}$ and $l_{2}$ as well as the $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ are correlated with each other, and therefore, measuring each of these four quantities are sufficient for the purpose of discriminating between tidal dissipation scenarios. We perform our analysis based on both the derivation in this paper (Equations 18) and that in Beuthe (2019). We find that the difference in the results is less than 5%. This close approximation indicates the applicability of (Equations 18) in capturing the tidal dissipation in the shell without a direct requirement of a priori constraint on the shell’s rheology.
Current Knowledge $k_{2}$ $\max\Delta R$ (cm) $\max\Delta E$ (cm) $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ (deg)
All studies N/A(62) N/A(28) 4 (11)
Park et al. (2024) 1 (87) 0.5 (40) 15 (20)
Park et al. (2024) 0.002 82 (134) 36 (61) 16 (21)
Park et al. (2024) 0.0002 116 (150) 51 (68) 22 (24)
Table 5: Range of the required precision for measuring $h_{2}$, $l_{2}$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$ for different scenarios of available and future measurements. “Current knowledge” implies the use of $\gamma_{I}$, $J_{3}$, and heat loss from the referenced studies. “All studies” refers to the use of estimates covering Howett et al. (2011), Thomas et al. (2016), Nadezhdina et al. (2016) , Hemingway & Mittal (2019), and Park et al. (2024). In the two bottom cases, measurement uncertainties on $k_{2}$ are assumed and other requirements are computed correspondingly. The requirements are given for obtaining 2-$\sigma$ confidence on the dominating region of the heat dissipation (with those associated with the 1-$\sigma$ confidence provided in parenthesis). N/A indicates that given the corresponding uncertainties, obtaining the 2-$\sigma$ confidence is not possible.
Figure 9a: Exploration of observables constraints from $J_{3}$, moment of inertia, libration amplitude, and heat loss rate based on Park et al. (2024). We presented PDFs of predictions for $k_{2}$, $h_{2}$, $l_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$. Both scenarios of tidal dissipation are considered here. The plots of the PDFs for the $k_{2}$ amplitude and phase lag are notably different than those of the $h_{2}$ and $l_{2}$ amplitudes and phases. Unit conventions are the same as in Figure 4.
Figure 9b: Exploration of the correlation between the model parameters associated with Figure 9a. Shear moduli are presented in GPa and other parameters are presented in their SI units. The ranges of the plausible interior parameters are shown in Table 4. Ice shell thickness, and shear moduli is plotted in logarithmic axes.
Figure 10: Panels (a)-(f): Tidal Love numbers and phase lags for heat dissipation in the shell and in the core based on the constraints by Park et al. (2024). Panels (g)-(j): Love numbers and phase lags for heat dissipation in the shell and in the core assuming that the real and imaginary parts of $k_{2}$ are measured within 0.002.
6.2 Constraining the interior structure with future observations
We use the favored values of the observables obtained in Section 6.1 along with the uncertainty of the measurements expected from future missions to constrain the main interior structure parameters. Due to the previously highlighted non-uniqueness of the solutions relying solely on the measurement of $k_{2}$, we perform two separate exemplar inversions associated with heat generation in the ice shell or in the core. The central value of the observation and the uncertainties adopted in the inversion are shown in Table 3. We use the favored values given in this table for $J_{3}$, contrary to the mean value of 181.066$\times$10-6 (adjusted for a reference radius of 251.99 km used here as compared to Park et al. (2024)). Our preferred value $J_{3}$ is within the 2-$\sigma$ uncertainty range of that suggested by Park et al. (2024). We use the lowest necessary precision for measuring $h_{2}$ to discriminate between the two scenarios of heat dissipation derived in Section 6.1. We conduct the inversion using MCMC sampling and the synthetic observations for two scenarios of measurement uncertainty given in Table 3 and discussed in Section 6.4. Each of our MCMC simulations includes approximately 600,000 accepted models which ensures that the parameter space is appropriately explored (algorithm details in Appendix C). The posterior distributions of the model parameters are demonstrated in Figures 11 and 12 for the two sets of inversions (only shown for the higher uncertainty case). Results based on both cases of higher and lower uncertainties are summarize in Table 6.
For the case of lower measurement precisions in Table 3, the results show that the suite of observations can constrain the thickness of the shell and the radius of the rocky core to within $\pm$2.7 km, and $\pm$7 km, respectively (3-$\sigma$ confidence level). When tighter constraints on the observables are adopted, the mean thickness of the shell can be constrained within approximately 1.2 km. The ocean density is constrained within $\pm$60 kg/m3 and $\pm$45 kg/m3 in the cases where the tidal dissipation is dominated in the core and the shell, respectively. This low precision is roughly equivalent to a knowledge of seawater salinity to within 40 psu (Millero et al., 2008, Earth’s ocean salinity is 35 psu or parts per thousand). An improved resolution of approximately $\pm$20 kg/m3 can be obtained if the higher precision scenario mentioned in Table 3 is used. The corresponding 20 psu salinity precision would allow for distinguishing the nominal 10 psu salinity of Postberg et al. (2009b) from an earthlike or greater salinity.
The viscosity and rigidity of the core are not well constrained if tidal dissipation is dominated in the shell because the range of $\eta_{\rm core}$ precludes a notable effect on the global tidal response of the core. This finding is consistent with the results of Ermakov et al. (2021) and Genova et al. (2024). The density of the core is constrained within 85 kg/m3 and 75 kg/m3 in case tidal dissipation occurs in the core and in the shell, respectively. If higher-precision measurements are acquired, the density of the core can be constrained within approximately 50 kg/m3.
In the case of a tidally active shell, the viscosity of the ice shell is constrained in the range of 4.4$\times$1013–9.4$\times$1013 Pa.s to satisfy the tidal heating constraint. In this tidal dissipation scenario, the shear modulus of the shell is constrained by the tidal Love numbers to be within 2.3-3.4 GPa. When tidal dissipation is dominated in the core, the shell rheology remains largely unconstrained, due to the insensitivity of the tidal heating to ice rheology in this scenario. We constrain the effective viscosity of the core in this scenario to be 9.1$\times$1011–1.4$\times$1012 Pa.s. A tighter constraint on the core viscosity is obtained (9.5$\times$1011–1.2$\times$1012 Pa.s), if the higher precision on the measurements are employed.In contrast to the viscosity of the core, its rigidity remains largely unconstrained.
Figure 11: Results of the MCMC inversion for the case where the tidal dissipation is dominantly in the core, provided the uncertainty on the measurements as shown in Table 3. The shear moduli are plotted in log axes. Results are presented for two sets of uncertainties presented Table 3.
Figure 12: Results of the MCMC inversion for the case where the tidal dissipation is dominantly in the ice shell and adopting the measurement uncertainty given in Table 3. The shear moduli are plotted in log axes. Results are presented for two sets of uncertainties presented Table 3.
Quantity Unit Tidal activity in the core Tidal activity in the shell
$D_{\rm shell}$ km 26.1–31.5 (26.1–28) 26–30 (26–28.4)
$\eta_{\rm shell}$ Pa.s 1.1$\times$1014–1020 (1015–1020) [4.4–9.4]$\times$1013 ([5–8.3]$\times$1013)
$\mu_{\rm shell}$ GPa 2.25–4 (2.65–4) 2.3–3.4 (2.65–3.1)
$\rho_{\rm shell}$ kg/m3 800–950 (900–950) 810–950 (860–950)
$\rho_{\rm ocean}$ kg/m3 1025–1145 (1020–1060) 1030–1120 (1040–1080)
$R_{\rm core}$ km 192–207 (193–200) 193–205 (193–202)
$\mu_{\rm core}$ GPa 1-65 (1–65) 2–70 (5–70)
$\rho_{\rm core}$ kg/m3 2190–2360 (2260–2355) 2210–2360 (2240–2355)
$\eta_{\rm core}$ Pa.s 9.1$\times$ 1011–1.4$\times$1012 (9.5$\times$1011–1.2$\times$1012) 1016–1020 (1016–1020)
Table 6: Posterior ranges of the model parameters from two exemplars of MCMC inversions using synthetic future measurements with expected accuracies presented in Table 3. Ranges are presented for two scenarios of measurement uncertainties.
6.3 Structural Heterogeneity in the ice shell
We assess the effect of lateral variations in ice shell thickness on the tidal response and its effect on the required measurements of $k_{2m}$ and $h_{2m}$. We build on the analysis of Berne et al. (2023) and conduct simulations of tidal deformation for four classes of models: 1. without lateral variations in ice shell thickness; 2. with lateral variations in ice shell thickness; 3. with a dissipative core and no lateral variations in ice shell thickness; and 4. with a dissipative core and lateral variations in ice shell thickness. For each model class, we vary the mean ice shell thickness across 15–50 km consistent with the finding based on the available measurements discussed in Section 6.1. We assume a core of viscosity $10^{20}$ Pa.s for the case of a non-dissipative core, and $10^{12}$ Pa.s for the case of a dissipative core. All other parameters are chosen to be the same as those used in Figures 1 and 2. To compute $k_{2m}$ and $h_{2m}$, we use a numerical finite element model (Berne et al., 2023). To generate crustal thickness variations, we apply the topography model in Park et al. (2024) (with minor modifications) evaluated up to $l=8$ to the outer surface of geometries and compensate non-hydrostatic components of this structure at the ice-ocean boundary based on the formulation for Airy-type isostasy described in Hemingway & Matsuyama (2017). The shape model suggests significant variations in the thickness of the shell (Figure 14), ranging between 5 and 50 km.
[…]
Figure 13: Top: gravitational Love numbers, $k_{20}$ (left column) and $k_{22}$ (right column), against mean ice shell thickness for models that assume a spherically symmetric ice shell (dotted line), an ice shell with lateral variations in thickness (solid line), or a dissipative core (dotted and solid red lines). Bottom: Similar to top row except showing displacement Love numbers $h_{20}$ and $h_{22}$ plotted against mean ice shell thickness. x-axes are plotted on a log-scale.
We also assess the effect of the ice shell’s structural heterogeneity on the maximum radial and horizontal tidal deformations on the surface compared to those derived in Section 6.1. The overall maximum surface deformation can differ in comparison to the spherically symmetric case used in previous studies. We compute the maximum displacements on the radial, northerly, and easterly directions for an ice shell with variable thicknesses and a spherically symmetric ice shell of the same average thickness equal to 30 km (Figure 15). In Figure 15, it is assumed that the tidal dissipation occurs in the core, and the viscosity of the ice shell is sufficiently high to assume a purely elastic behavior. The ice shell with heterogeneous thickness demonstrates variations in the radial and horizontal directions reaching 50 cm compared to a spherically symmetric shell (Figure 15). Larger deviations are observed for all components around the South pole, as expected, due to the significant variations in the thickness of the shell. The overall maximum radial displacement for the case of 3D structure is approximately 20 cm higher than the spherically symmetric case. The significant deviation of surface deformation demonstrated in these plots implies that for Enceladus, $h_{2}$ and $l_{2}$ can provide rough estimates of the actual surface displacements as commonly used for spherically symmetric bodies, and for a precise calculation of the displacement at any point on the surface the effect of the 3D structure is important. Since the globally maximum radial and horizontal displacements for a spherically symmetric body are smaller than those for a shell with 3D structural variations, the required accuracies presented in Section 6.1 derived by assuming spherical symmetry are conservative and thus remain valid.
Figure 14: Variation of Enceladus’s ice shell thickness based on the shape model by Park et al. (2024) assuming isostasy. Contours denote intervals of 2 km.
Figure 15: Amplitude of tidal deformation in radial, northerly, and easterly directions for two cases of an ice shell with variable thickness as shown in Figure 14 and a spherically symmetric shell with the same average thickness of 30 km. Here, it is assumed that the tidal dissipation is happening in the rocky core and the viscosity of ice shell is 1020 Pa.s, implying elastic deformations in the shell. Contours denote intervals of 0.05 m
6.4 Feasibility of the required measurements
The static and time-varying gravity field of Enceladus can be recovered by tracking the motion of a spacecraft (Park et al., 2016, 2020, 2024). The shape and topography of Enceladus can be derived using spacecraft measurements such as imaging (Park et al., 2015, 2019, 2020) or altimetry (Smith et al., 2001). The accuracy of recovered parameters typically depends on the measurement sensitivity, orbiting geometry, and spacecraft operating conditions. With reasonable assumptions, we can compute the expected precision of recovered Enceladus gravity and tidal parameters calculated in Section 6.1 presented in Table 3 (Smith et al., 2001; Park et al., 2015, 2016, 2019, 2020, 2024).
Previously, Park et al. (2020) simulated a scenario where a spacecraft is in a stable orbit around Enceladus with approximately 60 inclination and periapsis and apoapsis altitudes of 150 km and 200 km, respectively. Assuming a typical Deep Space Network (DSN) X-band tracking capability, $k_{2}$ and $J_{3}$ could be easily recovered with accuracies of $<0.002$ and $0.4\times 10^{-6}$ (Park et al., 2020), respectively. Considering a similar orbit geometry, but with a reasonable mission duration and a more capable X-up/Ka-down tracking system, $k_{2}$ could be recovered with an accuracy of $<0.0002$ and $0.04\times 10^{-6}$ (Park et al., 2020), respectively.
If an imaging system with high resolution and accurate pointing capabilities is used, temporal changes in surface features can be measured with an accuracy better than approximately half a pixel using a stereo reconstruction technique (Park et al., 2020). Assuming an imaging system with 1-meter resolution and $<1$-meter pointing accuracy, the libration amplitude can be measured with $<1$-meter accuracy. Furthermore, tidal deformation (both radial and horizontal) can be measured with an accuracy $<50$ cm. This scenario assumes images of the same surface feature are collected under varying observation geometries and at different phases of Enceladus relative to Saturn.
An onboard altimeter would be ideal for directly measuring radial deformation (e.g., Steinbrügge et al., 2015; Thor et al., 2020). By assuming a similar orbiting geometry and sampling strategy with an altimetry precision of 1 m, the radial deformation could be recovered with an accuracy $<20$ cm using surface tie points (i.e. repeat observation of the same surface point at different orbital phase).
7 Conclusions and discussion
Determining the interior properties of Enceladus remains challenging given the currently available observations. We explored the possibility of determining the internal structure of Enceladus and the tidal heating processes by means of a suite of geodetic measurements. We presented constraints on the interior structure of Enceladus by revisiting the latest analysis of geodetic data obtained by Cassini. Based on these analysis, we provide estimates of the plausible ranges of the currently unavailable tidal response, i.e., $k_{2}$, $h_{2}$, and $l_{2}$, and their associated phase lags. The tidal Love numbers of Enceladus compatible with the currently available observations span a wide range depending on the interior properties. We showed that the potential measurement of the $k_{2}$ and phase lag can be explained by tidal heat generated in the shell or in the rocky core, presenting a case of degeneracy that requires further constraints to be resolved. Therefore, neglecting the part of the parameter space associated with core-dominated tidal dissipation can result in an erroneous interpretation of the time-variable gravity field for the interior structure. Our analysis shows that measuring tidal displacements and/or phase lags can discriminate between core-dominated and shell-dominated tidal dissipation cases. If $k_{2}$ is measured within 0.002, the measurement of the maximum vertical and horizontal displacements within 82 cm and 40 cm are sufficient to obtain a 2-$\sigma$ confidence for discriminating between the two heat generation scenarios. The phase lag associated with $h_{2}$ differs significantly between the two cases of tidal dissipation. In the case of a core-dominated tidal dissipation and an elastic ice shell, $\epsilon_{h_{2}}$ is negligible, while for a shell-dominated tidal dissipation, $\epsilon_{h_{2}}$ can reach tens of degrees (10–50). The requirements for measuring $h_{2}$ and $l_{2}$ are relaxed when improved constraints on $k_{2}$ are obtained. If $k_{2}$ is constrained within 0.0002, measuring the vertical and/or tangential tidal displacements within 116 cm and 51 cm, respectively, is sufficient to discriminate between the two classes of tidal dissipation. We note that while measuring the tidal phase lags is probably necessary to constrain the heat generation mechanisms in the case of Europa and Ganymede as considered by Hussmann et al. (2016), here we demonstrate that in the case of Enceladus, the anelastic deformations can be large enough to constrain the first-order heat generation scenario. Meanwhile, measurements of phase lags for Enceladus can considerably improve our constraints on the rheological properties of the interior.
We note that the computed values for shell and core viscosity should be considered as the “effective” viscosity. The viscosity profiles can depend on the variation of parameters such as porosity, temperature, and the water saturation percentile with depth. Nevertheless, even if other mechanisms than viscoelastic dissipation are producing the inferred tidal heating, the signature on the tidal response will remain the same.
We also investigated the effect of the structural heterogeneities of the ice shell on the tidal response. Our results show that in the case of a largely elastic shell the surface displacements can significantly deviate from those of a spherically symmetric shell. These deviations can reach more than $\sim$50 cm for the thinnest parts of a heterogeneous shell compared to a spherically symmetric shell with the same global mean thickness. This observation is important for mission concept design purposes and for determining the locations at which the maximum displacements are measurable. We quantified the effect of a tidally active core on the tidal Love numbers of Enceladus with a heterogeneous ice shell. We found that if the core is tidally active to the extent needed to explain the steady-state heat budget, its effect on the observed $k_{2}$ in the presence of a heterogeneous shell can also be significant.
We conducted inversions using synthetic observables and demonstrated their efficacy for probing the main interior structure parameters such as the rheology, density, and thickness of the layers, using a feasible set of measurements and uncertainties. As one of the important factors of determining habitability, we demonstrated that high- and low-density ocean scenarios differing by more than 40 psu (see Postberg et al., 2009b; Glein et al., 2018; Glein & Waite, 2020; Fifer et al., 2022; Postberg et al., 2023) can be distinguished using geodetic measurements. Such different interpretations originate mainly from the underlying assumptions on the provenance of the plume material, the potential change in their composition during their transport from the ocean, and the undetermined extent of overturning circulation and possible ice-ocean interactions (Zhu et al., 2017; Soderlund, 2019; Kang & Flierl, 2020; Lobo et al., 2021; Kang et al., 2022; Kang, 2022; Meyer et al., 2025; Ames et al., 2025; Bouffard et al., 2025). This obtained constraints on the density is only marginal, but it is strengthened when tighter constraints from the gravity field are obtained. Constraining core density and rheology can help distinguish between a low-density porous core and a higher-density rocky core and help us understand its composition, formation, and evolution.
By extending the framework presented in this study to incorporate the future observations of total heat loss through thermal IR or microwave radiometry and comparison with the total tidally produced heat, we can understand whether Enceladus is in a thermal and orbital steady state, or alternatively, is experiencing a transient period during which the ocean may be becoming thicker or thinner. Such transient periods could occur in a cycle of freezing and melting due to a dynamically unstable orbit (Meyer & Wisdom, 2008) or in the case of an ongoing feedback between the moon’s thermal and orbital evolution of the body in the mean motion resonance (Ojakangas & Stevenson, 1986; Renaud et al., 2021). A transient ocean could also be the result of recent catastrophic events that affect the Saturnian system involving Enceladus (e.g., Asphaug & Reufer, 2013; Wisdom et al., 2022; Teodoro et al., 2023). By comparing the tidally generated heat in Enceladus with an independently measured surface heat loss rate, we can determine Enceladus’s thermal state and determine upper and lower bounds on the age of the subsurface ocean. The suite of needed geodetic measurements can be achieved by a future mission that provides additional constraints on the static gravity field, libration, tidal Love numbers, and rate of surface heat loss.
8 Acknowledgments
A. Bagheri acknowledges the support from the Swiss National Science Foundation by grant number P500PT_214435. A. Berne is supported by the Future Investigators in NASA Earth and Space Science and Technology (FINESST) Program (80NSSC22K1318).
Appendix A Viscoelastic models
Different models are proposed to capture the viscoelastic behavior of materials such as Maxwell, Andrade, Extended Burgers, and Sundberg-Cooper. The models differ in their representation of the transient regime between largely elastic to viscous dominated dissipation. If the frequency of excitations is close to the nominal Maxwell relaxation time of the body ($\tau_{M}=\eta/\mu$) or some layers of the body (i.e. $\tau_{M}(core)=\eta_{core}/\mu_{core}$ or $\tau_{M}(shell)=\eta_{shell}/\mu_{shell}$), the effect on the predicted response from different viscoelastic models can differ. However, if the excitation frequency (tidal period) is not close to the Maxwell relaxation time (e.g., in the case of Titan with a tidal period of $\sim$16 days), the choice of viscoelastic models does not change the result. For Enceladus, the orbital frequency is relatively close to the Maxwell relaxation time of the ice with its mechanical properties observed under Earth conditions. Therefore, Maxwell’s model may not be appropriate for predicting Enceladus’s tidal response.
Because periodic tidal deformations are small, we assume that the stress‐strain relationship is linear, such that the response of the material to shear forcing is described in the time domain by the creep function $J(t)$. The creep function for Maxwell’s model is written as:
$J(t)=J_{U}+\frac{t}{\eta},$ (A1)
whereas for the Andrade viscoelastic model, the creep function is given by (Andrade, 1962):
$J(t)=J_{U}+\beta t^{\alpha}+\frac{t}{\eta},$ (A2)
where $\alpha$ and $\beta$ describe a time-scale dependence of the anelastic relaxation strength. The associated dynamic compliance, which represents the response to a sinusoidally time-varying stress, has real and imaginary parts given by the following functions (Findley & Davis, 2013):
$\Re{(\hat{J}(\omega))}=J_{U}+\beta\Gamma(1+n)\omega^{-n}\cos(n\pi/2),$ (A3)
$\Im{(\hat{J}(\omega))}=-\beta\Gamma(1+n)\omega^{-n}\sin(n\pi/2)-1/\eta\omega,$ (A4)
where $\Gamma(1+n)$ is the Gamma function and $\omega$ is the angular frequency. Here, $\alpha$ (frequency dependence) and $\beta$ are empirical parameters that are not well constrained. $\alpha$ is usually assumed to be in the range 0.1–0.4 based on laboratory experiments (e.g., Jackson & Faul, 2010) and $\beta$ is suggested to be approximately $\mu^{\alpha-1}/\eta^{\alpha}$ (Castillo-Rogez et al., 2011). Currently, these parameters are not well constrained for most planetary bodies (see e.g. Nimmo et al., 2012; Nimmo & Faul, 2013; Bagheri et al., 2019; Petricca et al., 2025). Given the current lack of such constraints for Enceladus, we follow Andrade (1962) and Faul & Jackson (2015), and use $\alpha=0.2$ and $\beta=1$. We investigate the effect of variation in the Andrade parameters in Figure 17. The approximation for $\beta\approx\mu^{\alpha-1}/\eta^{\alpha}$ (Castillo-Rogez et al., 2011) is employed. We find that the change in the real parts of the Love number as a function of the changes in the Andrade parameters within the range of conventionally assumed end-member values remains less than 10% if the dissipation is dominated in the shell and less than 20% if the dissipation is dominated in the core.
The expressions for Maxwell’s model can be obtained by setting $\beta=0$ in the above equations. The relaxed (effective) shear modulus is thus calculable as (Findley & Davis, 2013):
$\mu=[J_{1}^{2}(\omega)+J_{2}^{2}(\omega)]^{-1/2},$ (A5)
and shear dissipation is
$Q^{-1}_{B}=J_{2}(\omega)/J_{1}(\omega).$ (A6)
Figure 16: Difference between different tidal observables as due to the use of the Maxwell and Andrade rheology: (a) tidal heating ($\Delta\dot{E}$), (b) maximum radial displacements ($\Delta R$), $k_{2}$ (c) and $h_{2}$ (d) Love numbers ($\Delta k_{2}$ and $\Delta h_{2}$) and associated phase lags ($\Delta\epsilon_{k_{2}}$ and $\Delta\epsilon_{h_{2}}$) (e, f). The choice of the viscoelastic model between Andrade and Maxwell’s models is typically more important as the tidal forcing periods become closer to the intrinsic relaxation time of the body. The choice of viscoelastic models for Enceladus can be potentially important because the orbital frequency of Enceladus (32.9 hr) is close to the estimated relaxation time for ice ($\tau_{M}=\eta_{ice}/\mu_{ice}$). The four narrow regions constrained by the red lines in panels (a) and (b) indicate the estimated tidal heat and the associated core viscosities (see Figure 1).
We show the differences between the tidal response parameters calculated based on Maxwell and Andrade viscoelastic models (for both the ice shell and the core), including the amplitudes of $k_{2}$ and $h_{2}$, their phase lags, and the resulting tidal heating (Figure 16). $h_{2}$ is relatively insensitive to the choice of the viscoelastic model for the rocky core due to the isolating effect of the ocean. In contrast, $k_{2}$ is sensitive to the choice of the viscoelastic model both for the ice shell and, to some extent, for the rocky core. The phase lag of $k_{2}$ can differ by as much as $\sim$14 between the Maxwell model and the Andrade model.
Figure 17: Effect of changing the Andrade parameter $\alpha$ on the tidal response of Enceladus. Right (Left) panel, for a model of tidal dissipation dominantly in the shell (core).
Appendix B Derivation of tidal heating partitioning between the shell and the core
We can compute the tidal heating in the specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body as a result of tidal perturbations. Note that Equation (5) cannot be used for the core or the surrounding layers simply by replacing $k_{2}/Q$ or $\Im(k_{2})$ of the whole body by the tidal Love numbers of the core or the layers computed independently of the rest of the body. Here, we use equation 37 in Tobie et al. (2005) to derive the contribution of the core to the global tidal heating:
$H_{tidal}(r)=-\frac{21}{10}\frac{\omega^{5}e^{2}R_{E}^{4}}{r^{2}}H_{\mu}\Im{(\mu(r))},$ (B1)
where $H_{tidal}$ is the radial distribution of the dissipation rate per unit volume, $r$ is the radius from the center, $H_{\mu}$ is the sensitivity kernel for shear dissipation given in Equation (33) in Tobie et al. (2005), and $\Im{(\mu(r))}$ is the imaginary part of the complex shear modulus which is assumed here to be constant throughout each layer. By integrating this expression from the center of the satellite to the core-ocean boundary, we obtain:
[…]
The integral terms can be written in terms of the imaginary part of the piecewise tidal Love number of the core following Equation 35 in Tobie et al. (2005):
[…]
Here, $\hat{k}_{2}^{core}$ is the piecewise gravitational tidal Love number associated with the core. For a homogeneous core, it can be computed from the deformation at the top of the core:
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}}$ (B4)
The tidal heating in the shell is computed using subtracting the contribution of the shell from the core:
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (B5)
Tidal heating in the shell can be computed separately with a similar approach:
[…]
$\hskip 26.46085pt-\frac{21}{2G}\omega^{5}e^{2}R_{E}^{4}\bigg{(}R_{E}\Im{(k_{2})}-R_{core}\Im{(\hat{k_{2}}^{core})}\bigg{)}.$ (B6)
Equations (B2), (B5), and (B) provide a method to calculate the partitioning of global tidal heating for a given interior structure model. These equations require a priori information, assumptions, or independent constraints on the rheological and structural properties of the core to compute the shear dissipation in the core.
This method can be used in a forward model or in the context of an inverse problem based on a sampling approach. However, we consider requirements on the geophysical measurement that do not directly depend on the interior model parameters. Thus, we extend our analysis to derive the tidal heating on the basis of potentially observable quantities. Quantifying the partitioning of the heat in the icy body is equal to calculating the contribution of the core and the shell to the imaginary part of the gravitational Love number $\Im{(k_{2})}$, as indicated from Equation (5). Beuthe (2019) uses membrane mechanics to obtain the contribution of each layer within the body to the imaginary part of the tidal Love number:
$\Im{(k_{2})}=\left|\frac{k_{2}+1}{k_{2}^{o}+1}\right|^{2}\Im{(k_{2}^{o})}-\xi_{2}|h_{2}|^{2}\Im{(\Lambda_{T})},$ (B7)
where $\xi_{2}=3\rho_{o}/5\rho_{b}$, $\rho_{o}$ and $\rho_{b}$ are the densities of the ocean and the whole body, $k_{2}$ and $h_{2}$ are the tidal Love numbers of the whole body, and $k_{2}^{0}$ is the fluid-crust gravitational Love number (assuming the crust is fluid), and $\Lambda_{T}$ is the shell spring constant introduced by Beuthe (2019). The contribution of the shell and the core to the total tidal heating can then be written as:
$\dot{E}_{\text{core}}=-\frac{5}{2}\frac{\omega R}{G}\left|\frac{k_{2}+1}{k_{2}^{0}+1}\right|^{2}\Im(k_{2}^{0})\langle|U_{2}^{(T)}|^{2}\rangle,$ (B8)
$\dot{E}_{\text{shell}}=\frac{5}{2}\frac{\omega R}{G}\xi_{2}\,|h_{2}|^{2}\Im(\Lambda_{T})\,\langle|U_{2}^{(T)}|^{2}\rangle,$ (B9)
where,
$\langle|U_{2}^{(T)}|^{2}\rangle=(\frac{21}{5})(\omega R)^{4}e^{2}.$ (B10)
The fluid crust gravitational Love number can be computed from Equation D.2 in Appendix D and Table 9 in Beuthe (2019). Once the tidal heating rate in the shell is computed, the partition of the heat produced in the core can be computed by subtracting the total heat from that of the shell. Following Beuthe (2015), the term $\Lambda_{T}$ in Equation (B8) is the complex spring constant of the shell and is computed as:
$\Lambda_{T}=\Lambda_{M}+\Lambda_{\rho}+\Lambda_{\chi}+\Lambda_{\omega}$ (B11)
The term $\Lambda_{M}$ refers to the contributions by the membrane spring constant that has the dominant effect, and $\Lambda_{\rho}$, $\Lambda_{\chi}$, and $\Lambda_{\omega}$ are corrections due to the density contrast between the shell and the ocean, compressibility correction, and dynamical correction, respectively. Here, we consider the effect of the two leading terms, $\Lambda_{M}$ and $\Lambda_{\rho}$, the latter two terms having a negligible contribution (Beuthe, 2015). The contributions of the membrane spring estimate and density contrast are (Beuthe, 2015):
$\Lambda=f_{\mu}\hat{\mu}\epsilon,$ (B12)
$\Lambda_{\rho}=f_{\rho}\frac{\delta\rho}{\rho}\epsilon.$ (B13)
Here, $\delta\rho$ is crust-ocean density contrast, $\epsilon=d/R$ and $f_{\mu}$ and $f_{\rho}$ are:
$f_{\mu}=\frac{8(1+{\nu})}{(5+{\nu})},$ (B14)
$f_{\rho}=\frac{1+5{\nu}}{5+{\nu}},$ (B15)
and,
$\hat{\mu}=\frac{\mu}{\rho gR},$ (B16)
where $\mu$ is the shear modulus, $\rho$ is the density, and $\nu$ is the Poisson’s ratio of the shell. $\nu$ can be written in terms of the shear and bulk moduli:
$\nu=\frac{3K-2\mu}{6K+2\mu}.$ (B17)
This formulation provides a means for obtaining the tidal heating in the shell by measuring the tidal Love number $h_{2}$ (real and imaginary parts) and independent constraint on the imaginary part of the complex shear modulus of the shell, $\Lambda_{T}$, as shown in Equation (B8). Here, we extend the formulation to develop expressions for the partitioning of heat between the shell and the core, which do not require a priori knowledge of the shell rheology. Equation 125 in Beuthe (2015) provides an expression between the displacement and potential Love numbers, the rigidity of the shell, and the density structure of the body:
$k_{2}+1=(1+\Lambda_{T})h_{2}-5\frac{\delta\rho}{\rho}\epsilon.$ (B18)
The rigidity constant of the shell can be written in terms of the other parameters as:
$\Im{(\Lambda_{T})}=\Im\bigg{(}\frac{1}{h_{2}}{(k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon)}\bigg{)},$ (B19)
Combining Equations (B8) and (B19), we can rewrite the expression of tidal heating in the shell as:
$\dot{E}_{\text{shell}}=\frac{105}{10}\frac{(\omega R)^{5}}{G}\xi_{2}\,|h_{2}|^{2}\bigg{[}\Im\bigg{(}{\frac{1}{h_{2}}\big{(}k_{2}+1+5\frac{\delta\rho}{\rho}\epsilon\big{)}}\bigg{)}\bigg{]}\,e^{2}.$ (B20)
This formulation provides the tidal heating in the shell by measurements of the real and imaginary parts of the tidal Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers of the body can be constrained using the static gravity field and libration, and are easier to constrain compared to the rheology of the layers as required in Equation (B8). Tidal heating in the core can be computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for all other icy satellites for which the necessary geodetic observations are or will become available.
Appendix C Markov Chain Monte Carlo Internal Structure Inversion
We adopt the Metropolis algorithm to sample the posterior distribution in the model space (Mosegaard & Tarantola, 1995). This algorithm, which samples the model space in a random manner, ensures that models that fit the data better and are simultaneously consistent with prior information are sampled more frequently. The Metropolis algorithm samples the model space with a sampling density proportional to the posterior probability density and, therefore, ensures that low-probability areas are sampled less excessively. Assuming that data uncertainty is Gaussian distributed and that observational uncertainties and calculation errors among the data sets considered are independent, the likelihood function can be written as
$\mathcal{L}(\textbf{m})\propto\Pi\exp\bigg{(}-\frac{|d_{obs}^{i}-d_{cand}^{i}(\textbf{m})|^{2}}{2\sigma_{i}^{2}}\bigg{)}$ (C1)
where $i$ runs over the observables, m is the vector consisted of the model parameters, and $d_{obs}$ and $d_{cand}$ are observed and computed candidate data in the algorithm, respectively, and $\sigma_{i}$ is data uncertainty.
Appendix D Probability Distribution Function (PDF) of interior model parameters and correlation
Figure 18: Exploration of the key geophysical interior model parameters corresponding to with Figure 4. Multiple peaks observed in the PDFs of the viscosities of ice shell and the core correspond to separate tidal dissipation regimes that can match the inferred global heat loss. These peaks, and the correlation observed between the viscosities of the shell and rock correspond to the regions indicated in Figure 1. Shear moduli are presented in GPa and other parameters are presented in their SI units.
The sensitivity of the adopted set of measurements used in Figure 18 to parameters such as the density and rigidity of the layers is small. Multiple peaks are apparent in the PDFs of the viscosities of the core and the shell, which are associated with the multiple regions highlighted in Figure 1 that are consistent with the estimated heat loss. These distinct peaks are indicative of tidal dissipation dominated by viscous deformation versus anelastic deformation. In other words, these peaks correspond to viscosities that are lower than or higher than the viscosity associated with the maximum heat production rate.
When considering the constraints on the geodetic observables of Park et al. (2024) and separating the interior parameters for the two tidal dissipation scenarios, we obtain constraints on the interior parameters and correlations between them, as shown in Figures 21 and 22.
Figure 19: Exploration of the correlation between the model parameters associated with Figure 5. Shear moduli are presented in GPa and other parameters are presented in their SI units.
Figure 20: Exploration of the correlation between the model parameters associated with Figure 6. Shear moduli are presented in GPa and other parameters are presented in their SI units.
Figure 21: model parameters for the case where at least 90% of the tidal heating is focused in ice shell. The tidal heating in the shell is calculated using the formulation by (Beuthe, 2019) (Equation (B8)) (Supplementary material figure).
Figure 22: Model parameters for the case where at least 90% of the tidal heating is focused in the rocky core. The tidal heating in the shell is calculated using the formulation in (Beuthe, 2019) (Equation (B8)) and the heating in the core is calculated by subtracting the heat in the shell from the total tidal heating computed based on Equation (5) (Supplementary material figure).